An institute conducted a clinical trial of its methods for gender selection. The results showed that 526 of 921 babies born to parents using a specific​ gender-selection method were boys. Use the sign test and a 0.1 significance level to test the claim that the method increased the likelihood of having a boy.

Find the null and alternative hypothesis.

H0:

  • p<0.5
  • p>0.5
  • p=0.5
  • p≠0.5

H1:

  • p<0.5
  • p=0.5
  • p≠0.5
  • p>0.5

If we consider + to represent a boy, then how many of each sign is there?

Positive Signs:

Negative Signs:

Total Signs:

What is the p-value?  (Round to three decimal places.)

What is the conclusion about the null?

What is the conclusion about the claim?

Please explain how to find the p-value on a TI-84 calculator. Thanks.
Answer #1

Null hypothesis, H0:p=0.5

Alternative hypothesis, H1:p>0.5

Positive Signs: 526

Negative Signs: 921-526=395

Total Signs: 921

There is enough evidence to accept the claim that the method increased the likelihood of having a boy.

p-value on a TI-84 calculator:

Right Tailed z-test:

1) Calculate z_calc (z_test) (here it is z_calc=(526-460.5)/sqrt(230.25)=4.3166

2) 2nd DISTR

3) Scroll down to normalcdf(

4) ENTER

5) Now enter: z_calc, 1000, 0,1)

6) ENTER

7) Output is the P-value